OpenStudy (anonymous):

Can I see someone solve, simplify the expression: (b^2+5b+6) divided by (3bc) divided by (b^2-9) divided by (6bc)

OpenStudy (anonymous):

you'll get something like this first :\[6bc ( b^2 + 5b + 6)/ 3bc(b^2-9)\]

OpenStudy (anonymous):

then you can cancel the 6bc with 3bc and you'll get :\[2(b^2+5b+6)/(b^2-9) = 2(b+3)(b+2)/(b-3)(b+3) = 2(b+2)/(b-3)\]

OpenStudy (anonymous):

2(b^2 + 5b + 6) / (b^2 - 9) 2(b + 3)(b + 2) / (b+3)(b-3) 2(b+2) / (b-3) (2b + 4) / (b-3)

OpenStudy (anonymous):

2(b+2)/(b-3) , sorry lol, lagging

OpenStudy (anonymous):

I can just do it fast. Ego strong here. :D

OpenStudy (anonymous):

lol, I was lagging =P

OpenStudy (anonymous):

Got that too

OpenStudy (anonymous):

Though I commend you for using the equation. And no, you still can't do it faster than me. ^_^

OpenStudy (anonymous):

lol, it's not about how fast you solve it, it's about whether you did it correctly or not :) + whether you helped someone or not ^_^

OpenStudy (anonymous):

It helped!

OpenStudy (anonymous):

It's about beating you, sstarica, that's all it's about.

OpenStudy (anonymous):

DX

OpenStudy (anonymous):

SStarica said 2(b+2)/(b-3) and Supervisor said (2b+4)/(b-3) and in my study book it says the answer is 2(b+2) / (b-3) my hero.